OpenStudy (anonymous):

finding vertex in standard form? y=-x^2+2x-5

OpenStudy (anonymous):

I understand how to get the x coordinate but i keep getting 8 for the y coordinate, and the book says it is 1, 6. can someone help me find my error

OpenStudy (mayankdevnani):

\[\large \bf ax^2+bx+c=0\] and discriminant,\(\large \bf D=b^2-4ac\) to find vertex ,we use this formula :- \[\large \bf Vertex(\frac{-b}{2a},\frac{-D}{4a})\]

OpenStudy (mayankdevnani):

can you now solve it ? @bastiennecroix

OpenStudy (anonymous):

The method we learned for finding the y is just plugging the x back into the quadratic

OpenStudy (anonymous):

Is it possible to find it that way? plugging it back in ? @mayankdevnani

OpenStudy (mayankdevnani):

yeah.Just plug all the values in formula and definitely,you will get correct answer.

OpenStudy (anonymous):

so, (-1)^2 + 2(1) + 5 ?

OpenStudy (anonymous):

sorry, the original quadratic is y= -x^2+2x+5

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

i can tell you one answer and other answer is for you :)

OpenStudy (anonymous):

see im confused because when that's added up, it's 8.

OpenStudy (mayankdevnani):

to find x-coordinate, \[\large \bf \frac{-b}{2a}=\frac{-2}{2(-1)}=\frac{-2}{-2}=1\]

OpenStudy (mayankdevnani):

therefore,x-coordinate is 1

OpenStudy (anonymous):

Yes. and so to find the y, we need to plug the x back into the quadratic

OpenStudy (anonymous):

so: -(1)^2 + 2(1) +5

OpenStudy (mayankdevnani):

yeah !!

OpenStudy (anonymous):

and that equals 8, unless my brain is messing up

OpenStudy (anonymous):

So I was right? My textbook says the vertex is 1, 6

OpenStudy (anonymous):

I've been going through this problem forever trying to see where I went wrong

OpenStudy (mayankdevnani):

oh!!! \[\large \bf -(1)^2+2+5=-1+2+5=6\]

OpenStudy (anonymous):

I thought squared negatives were always positive

OpenStudy (mayankdevnani):

no,the negative sign is out of the bracket,so +ve squared is positive

OpenStudy (anonymous):

OH! thank you so much!!!!!!

OpenStudy (mayankdevnani):

welcome :)